Keller Cosmetics maintains an operating profit margin of 9.00% and a sales-to-assets ratio of 3.90. It has assets of $700,000 and equity of $500,000. Assume that interest payments are $50,000 and the tax rate is 30%.

 

a. What is the return on assets? (Enter your answer as a percent rounded to 2 decimal places.)


b. What is the return on equity? (Enter your answer as a percent rounded to 2 decimal places.)
Answer #1


Given Information

Operating Profit Margin = 9%

Sales - To Assets Ratio = 3.90

Assets = $700,000

Equity = $500,000

Interest Payments = $50,000

Tax Rate = 30%


Using the Formula of Sales to Asset Ratio

Sales to Assets Ratio \(=\frac{\text { Sales }}{\text { Assets }}\)

\(3.90=\frac{\text { Sales }}{\$ 700,000}\)

\(3.90 \times \$ 700,000=\) Sales

\(\$ 2,730,000=\) Sales

Using the Formula of Operating Profit Margin

$$ \begin{aligned} \text { Operating Profit Margin } &=\frac{\text { Operating Profit }}{\text { Net Sales }} \times 100 \\ 9 \% &=\frac{\text { Operating Profit }}{\$ 2,730,000} \\ \$ 245,700 &=\text { Operating Profit } \end{aligned} $$


Operating Profit = $245,700

Particulars Values 
Operating Profit (EBIT)$245,700
Less : Interest Payment ($50,000)
Profit before Taxes $195,700
Less : Taxes @30%($58,710)
Earnings After Taxes or Net Income $136,990



a.

Using the Formula of Return on Assets

Return on Assets \(=\frac{\text { Net Income }}{\text { Total Assets }}\)

Return on Assets \(=\frac{\$ 136,990}{\$ 700,000}\)

Return on Assets \(=0.1957\) or \(19.57 \%\)


b.

Using the Formula of Return on Equity

Return on Equity \(=\frac{\text { Net Income }}{\text { Shareholder's Equity }}\)

Return on Equity \(=\frac{\$ 136,990}{\$ 500,000}\)

Return on Equity \(=0.27398\) or \(27.40 \%\)

answered by: Allen